OpenStudy (anonymous):

For a wedding, Shereda bought several dozen roses and several dozen carnations. The roses cost $15 per dozen, and the carnations cost $8 per dozen. Shereda bought a total of 17 dozen flowers and paid a total of $192. How many roses did she buy?

OpenStudy (texaschic101):

15r + 8c = 192 r + c = 17 ---> c = 17 - r Now just sub 17 - r in for c in the second equation and solve for r (roses).

OpenStudy (anonymous):

Thank you this really helped out!